solve the given equations using exact differential

equations.

1. (6x + y2)dx + y(2x − 3y)dy = 0


2. (y2 − 2xy + 6x)dx − (x2 − 2xy + 2)dy = 0, when x = 1, y = 2


3. v(2uv2 − 3)du + (3u2v2 − 3u + 4v)dv = 0, when u = 1, v = 1


4. (1 + y2 + xy2)dx + (x2y + y + 2xy)dy = 0


5. (w3 + wz2 − z)dw + (z3 + w2z − w)dz = 0, when w = 4, z = 2


1
Expert's answer
2021-10-01T00:50:31-0400

"\displaystyle
(6x+y^2 )dx - (2xy - 3y^2)dy= 0\\
\text{Let $M = 6x+y^2 $ and $N = 2xy - 3y^2$}\\
\text{Let $F(x,y) = \int (2xy - 3y^2)dy$}\\
\implies F(x,y) = xy^2-y^3 + h(x)\\
\text{Differentiating and equating it to M, we have}\\
y^2 +h'(x) = y^2+6x\\
\implies h'(x) = 6x\\
h(x) = 3x^2\\
\therefore F(x,y) = xy^2-y^3 + 3x^2\\
\text{The solution to the exact differential equation is}\\
xy^2-y^3 + 3x^2=c\\
2.
(y^2-2xy+6x)dx - (x^2-2xy+2)dy= 0\\
\text{Let $M = y^2-2xy+6x $ and $N = -(x^2-2xy+2)$}\\
\text{Let $F(x,y) = \int -(x^2-2xy+2)dy$}\\
\implies F(x,y) = -x^2y+xy^2-2y + h(x)\\
\text{Differentiating and equating it to M, we have}\\
y^2 -2xy+h'(x) = y^2-2xy+6x\\
\implies h'(x) = 6x\\
h(x) = 3x^2\\
\therefore F(x,y) =-x^2y+xy^2-2y+ 3x^2\\
\text{The solution to the exact differential equation is}\\-x^2y+xy^2-2y+ 3x^2=c\\
\text{When x= 1, y=2, c = 1, therefore the particular solution is}\\
-x^2y+xy^2-2y+ 3x^2=-1\\
3.
v(2uv^2-3)du + (3u^2v^3-3u+4v)dv= 0\\
\text{Let $M = v(2uv^2-3) $ and $N = 3u^2v^2-3u+4v$}\\
\text{Let $F(u,v) = \int (2uv^3-3v)du$}\\
\implies F(u,v) = u^2v^3-3uv + h(v)\\
\text{Differentiating and equating it to M, we have}\\
3u^2v^2-3u+h'(x) = 3u^2v^2-3u+4v\\
\implies h'(v) = 4v\\
h(v) = 2v^2\\
\therefore F(u,v) =u^2v^3-3uv+ 2v^2\\
\text{The solution to the exact differential equation is}\\u^2v^3-3uv+ 2v^2=c\\
\text{When u = 1, v=1, c = 0, therefore the particular solution is}\\
u^2v^3-3uv+ 2v^2=0\\
4.
(1+y^2+xy^2)dx + (x^2y+y+2xy)dy= 0\\
\text{Let $M = 1+y^2+xy^2 $ and $N = x^2y+y+2xy$}\\
\text{Let $F(x,y) = \int x^2y+y+2xydy$}\\
\implies F(x,y) = \frac{x^2y^2}{2}+\frac{y^2}{2}+xy^2 + h(x)\\
\text{Differentiating and equating it to M, we have}\\
y^2+xy^2+h'(x) = 1+y^2+xy^2\\
\implies h'(x) = 1\\
h(x) = x\\
\therefore F(x,y) =\frac{x^2y^2}{2}+\frac{y^2}{2}+xy^2 +x\\
\text{The solution to the exact differential equation is}\\-x^2y+xy^2-2y+ 3x^2=c\\
5.

(w^3+wz^2-z)dw + (z^3+w^2z-w)dz= 0\\
\text{Let $M = w^3+wz^2-z $ and $N = z^3+w^2z-w$}\\
\text{Let $F(w,z) = \int w^3+wz^2-zdw$}\\
\implies F(w,z) = \frac{w^4}{4}+\frac{w^2z^2}{2}-wz+h(z)\\
\text{Differentiating and equating it to M, we have}\\
w^2z-w+h'(x) = z^3+w^2z-w\\
\implies h'(x) = z^3\\
h(z) = \frac{z^4}{4}\\
\therefore F(x,y) =\frac{w^4}{4}+\frac{w^2z^2}{2}-wz+\frac{z^4}{4}\\
\text{The solution to the exact differential equation is}\\\frac{w^4}{4}+\frac{w^2z^2}{2}-wz+\frac{z^4}{4}=c\\
\text{When w= 4, z=2, c = 92, therefore the particular solution is}\\
\frac{w^4}{4}+\frac{w^2z^2}{2}-wz+\frac{z^4}{4}=92"